At 8:00 A.M., a train leaves a station and travels at a rate of 48 mph. At 9:00 A.M., a second train leaves the same station on the same track and travels in the direction of the first train at a speed of 64 mph. At 10:00 A.M., how far apart are the two trains?
Answer by srinivas.g(540)

speed of first train = 48 mph
BY 10 AM ,it travels for 2 hrs
distance traveled = speed x time
=48*2
=96 miles
speed of second train = 64 mph
BY 10 AM , it travels for 1 hr
distance covered = 64*1
=64 miles
distance b/w these trains = 96-64
= 32 miles
result is 32 miles